OpenStudy (anonymous):

Suppose E(Z)=7 and E[Z(Z-1)]=200. Find V(Z).

OpenStudy (tkhunny):

Please define V(Z) in terms of expected values.

OpenStudy (anonymous):

V(Z)=E[(Z-u)^2]

OpenStudy (tkhunny):

Very good. Now expand that expression inside the Expected Value brackets.

OpenStudy (anonymous):

\[\sum_{}^{}(Z-u)^2p(z)\]

OpenStudy (tkhunny):

What is that? \(E\left[(Z-\mu)^{2}\right] = E\left[Z^{2} - 2Z\mu + \mu^{2}\right]\) That is what "expand inside the ... brackets" means. Now what?

OpenStudy (anonymous):

the summation is how you find an Expected value (E())

OpenStudy (anonymous):

sorry i had misread what you typed

OpenStudy (anonymous):

I don't know what to do from here

OpenStudy (tkhunny):

You should have rules for linear combinations of expected values: \(E\left[Z^{2} - 2Z\mu + \mu^{2}\right] = E\left[Z^{2}\right] -2\mu E\left[Z\right] + \mu^{2}\) Does that look plausible?

OpenStudy (anonymous):

yeah okay.

OpenStudy (anonymous):

so how do I find E[Z^2] and mu? I can replace E[Z] with the value given...

OpenStudy (tkhunny):

Given in the problem statement. Keep in mind that \(E[Z] = \mu\)

OpenStudy (anonymous):

duh... i'm an idiot. and E[Z^2]?

OpenStudy (tkhunny):

You are SO CLOSE!! E[Z]=7 E[Z(Z-1)]=200 E[Z(Z-1)] = E[Z^2 - Z] = E[Z^2] - E[Z] = E[Z^2] - 7 = 200

OpenStudy (anonymous):

ohhhh so 207=E[z^2]!

OpenStudy (anonymous):

then its just adding and multiplying!

OpenStudy (tkhunny):

That easy, once you see all the pieces!!

OpenStudy (anonymous):

Thanks so much!